OpenStudy (anonymous):

Find the sum and simplify and show all work and steps please..will give medals!! 2x/( x^2-x-12) + 3/ (x^2+4x+3)

OpenStudy (mertsj):

Factor both denominators for the first step

OpenStudy (mertsj):

What are the factors of x^2-x-12 ?

OpenStudy (anonymous):

I have no idea. Im very confused with this question

OpenStudy (mertsj):



OpenStudy (mertsj):

Think of two numbers whose product is -12

OpenStudy (anonymous):

6 & -2?? 12 & -1 ??

OpenStudy (mertsj):

When you add the two numbers, you have to get -1 so think of some more possibilities.

OpenStudy (anonymous):

4 & -3

OpenStudy (mertsj):

4+(-3)=1 The sum must be -1

OpenStudy (anonymous):

-4 & 3

OpenStudy (mertsj):

Good.

OpenStudy (mertsj):



OpenStudy (mertsj):

So now you have factored the first denominator. Try the second one.

OpenStudy (anonymous):

X^2+4x+3 = (x+3)(x+1)

OpenStudy (mertsj):

So now we have this problem: \[\frac{2x}{(x-4)(x+3)}+\frac{3}{(x+1)(x+3)}\]

OpenStudy (mertsj):

Now we have to get the common denominator so we can add the two fractions.

OpenStudy (mertsj):

For the first denominator we need the factors (x-4)(x+3) For the second denominator we need the factors (x+1)(x+3)

OpenStudy (mertsj):

So the factors we need are (x-4) and (x+3) and (x+1) and that is the common denominator. Are you with me so far?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

So the first fraction will have to be multiplied by (x+1)/(x+1) to get the denominator right.

OpenStudy (mertsj):

The second fraction will have to be multiplied by (x-4)(x-4) to get the denominator right. Do you agree?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

\[\frac{2x}{(x-4)(x+3)}\times \frac{(x+1)}{(x+1)} + \frac{3}{(x+1)(x+3)}\times \frac{(x-4)}{(x-4)}\]

OpenStudy (mertsj):

So now if we do the multiplication we have:

OpenStudy (mertsj):

\[\frac{2x^2+2x+3x-12}{(x+1)(x+3)(x-4)}\]

OpenStudy (mertsj):

Still ok with this?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

So let's add the like terms and finish up.

OpenStudy (mertsj):

\[\frac{2x^2+5x-12}{(x+1)(x+3)(x-4)}\]

OpenStudy (anonymous):

Is that the final answer?? It doesnt simplify down anymore??

OpenStudy (mertsj):

You could try to factor the numerator but I don't think it will factor. Are there factors of -24 whose sum is 5?

OpenStudy (anonymous):

I dont think so. The correct answer would be 2x^2+5x-12 / (x+1)(x+3)(x-4) ?

OpenStudy (mertsj):

Oh it will factor. The factors are (2x-3)(x+4)

OpenStudy (mertsj):

So now we have: \[\frac{(2x-3)(x+4)}{(x+1)(x+3)(x-4)}\]

OpenStudy (mertsj):

But there are no common factors to cancel so our answer is correct.

OpenStudy (anonymous):

So the (2x-3)(x+4) and so on the correct answerr??

OpenStudy (mertsj):

The answer is: \[\frac{2x^2+5x-12}{(x+1)(x+3)(x-4)} or \frac{(2x-3)(x+4)}{(x+1)(x+3)(x-4)}\]

OpenStudy (mertsj):

because they are two forms of the same answer

OpenStudy (mertsj):

So you can take your pick.

OpenStudy (anonymous):

Oh okay i understand (: thank you so much!

OpenStudy (mertsj):

yw